What is the critical angle for light traveling from crown glass into air?
1
Expert's answer
2020-11-25T07:12:58-0500

Let's write the Snell's Law:


"n_isin\theta_i=n_rsin\theta_r,"

here, "n_i=1.52" is the index of refraction of the incident medium (or the crown glass), "n_r=1.0" is the index of refraction of the refractive medium (or the air), "\theta_i" is the angle of incidence, "\theta_r" is the angle of refraction.

The critical angle is defined as the angle of incidence that provides an angle of refraction of "90^{\circ}." Therefore, we can write:


"n_isin\theta_c=n_rsin90^{\circ},""sin\theta_c=\dfrac{n_r}{n_i},""\theta_c=sin^{-1}(\dfrac{n_r}{n_i})=sin^{-1}(\dfrac{1.0}{1.52})=41.1^{\circ}."

Answer:

"\theta_c=41.1^{\circ}."